If it takes 100m to stop a car initially moving at 100 m/s what distance is required to stop a car moving at 200 m/s under the same condition?

Show work using energy equations only.
a. 50m
b. 100m
c. 200m
d. 400m
Answer #1

Using work energy equation

WOrk done = Change in kinetic energy

F*S = 0.5m*(v^2 - u^2)

F*100 = 0.5*m*(100^2 - 0 )

F/m = 50

when speed is 200 m/s

F*S = 0.5*m*(200^2 - 0)

S = 0.5*200^2 / 50

= 400 m

so the correct answer is option d)400 m